For a charging capacitor in an RC circuit with R = 100 kΩ and C = 1 µF, and in series with a battery with emf E = 12 V, (a) how long does it take for the capacitor to charge up to 80% of its final charge? (b) What is the current in the circuit at that point in time?
Answer #1

R = 100 kΩ and C = 1 µF
E = 12 V

For Charging RC Circuit,
q = Qf*( 1 - e^(-t/RC))
0.8*Q = Q * ( 1 - e^(-t/(100*1000 * 1 * 10^-6))
t = 0.161 s

(b)
Current in the cirucit,
I = Io * e^(-t/RC)
I = 12/(100*10^3) * e^(-0.161/(100*1000 * 1 * 10^-6))
I = 2.4 * 10^-5 A